Circular No. 5717

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993F IN ANONYMOUS GALAXY C. Pollas reports his discovery of an apparent supernova of mag ∼ 18.5

on two technical pan films (limiting mag 21-22) taken Jan. 18.97 and 20.94 UT by D. Albanese and himself with the 0.9-m Schmidt telescope at the Observatoire de la Côte d’Azur. The candidate, at α = 7^{h}54^{m}26^{s}.45, δ = +20^{o}13′04′′.8 (equinox 1950.0), is superimposed on the nuclear region of a galaxy of mag 18, being roughly 0′′.6 east and 1′′.5 south of the galaxy’s center. No such stellar image is present on a technical pan plate (similar limiting magnitude) obtained in 1990 January, though there is perhaps a condensation of mag ∼ 20 nearly at the position of the supernova visible on an old film. A nearby star (m_{pg} = 19) has end figures 20^{s}.92, 14′17′′.7.

4U 0614+09 = V1055 ORIONIS S. Brandt, N. Lund, and A. J. Castro-Tirado, Danish Space Research

Institute, report: “A new x-ray burst from the low-mass x-ray binary 4U 0614+09 has been detected by the WATCH all-sky monitor on the EU-RECA satellite. The burst started at Mar. 2.59408 UT and in 16 s reached a peak flux in excess of 20 Crabs in the energy band 6-15 keV. The burst was the second observed from this source within a two-week period and was ∼ 50 percent brighter than the burst seen on Feb. 17 (IAUC 5710). A 4σ excess flux was also detected above 15 keV.”

α ORIONIS

A. K. Dupree, E. Guinan, and M. Smith further report, in addition to the information on IAUC 5716: “Photometry at Villanova on Feb. 28 UT suggests that α Ori has paused in its decline in brightness at V ∼ +0.9. International Ultraviolet Explorer observations were acquired as a part of the ongoing monitoring program on Feb. 24. Radial velocity measurements in the Hα region continue at the McMath Solar Telescope at Kitt Peak.”

CORRIGENDA On IAUC 5678, ‘Supernova 1992bn in Anonymous Galaxy’, line 5, for

1′′.2 east and 6′′.8 north read 1′′.2 west and 6′′.8 south

On IAUC 5712, ‘Corrigenda’ for plates taken by M. J. Drinkwater read films taken by C. P. Cass

1993 March 3 Daniel W. E. Green